Jacinta plays basketball. She scores 5 goals in 3 games. If she continues to score goals in this pattern how many games will she need to play to score a total of 40 goals?
Answer by checkley77(12844)

YES 5/3 is a division.
5 points scored in 3 games=5/3=1.6667 points per game.
-------------------------------------------------------
5/3=40/X
5X=120
X=120/5
X=24 GAMES ARE NEEDED TO SCORE 40 GOALS AT HER PRESENT RATE.
PROOF:
5/3=40/24
120=120